# Radiation properties and images of loop quantum Reissner-Nordström black hole with a thin accretion disk

### v1

@@ Abstract
We investigate the characteristics of circular geodesics around loop quantum Reissner-Nordström black hole (LQRNBH) and the radiation properties and observational appearance of a thin accretion disk around it. By calculating the shadow radius and utilizing observational data from M87* and Sgr A*, we derive constraints on the quantum parameter $\zeta$ and charge parameter $Q$. The timelike circular geodesics around LQRNBH and the influence of the model parameters on the circular motion are also discussed. It is found that contrary to the case of the parameter $Q$, the increase of the quantum parameter $\zeta$ leads to the increase of the radius of the innermost stable circular orbit (ISCO). Then, by considering a thin accretion disk model, various radiation properties of the LQRNBH and the effects of the model parameters on them are studied. Concrete examples are provided for quantitative comparison of the radiation properties between LQRNBH and Schwarzschild black hole. With the ray-tracing method, the isoradial curves, redshift distributions, and the observed radiation fluxes of the direct and secondary images of the LQRNBH with the thin accretion disk for various model parameters and observation angles are numerically calculated and discussed. These results are beneficial for us to understand the physical consequences of loop quantum gravity effect.
@@ I Introduction
The LQBHs have been studied extensively from various aspects. The quasinormal modes of these spherical LQBHs have been explored by several recent works [27, 28, 29, 30, 31, 32, 33, 34]. The correspondence between grey-body factors and quasinormal modes of these LQBHs were also studied [35]. The features of the shadows of these quantum black holes were studied in [36, 37, 38, 39], and the constraints on the parameter encoding the quantum gravity effect from Event Horizon Telescope (EHT) observations [34, 40, 41] and gravitational wave observation [42] were also discussed. Furthermore, the optical appearance of the accretion disk around LQBHs and the strong gravitational lensing effects of the LQBHs were explored in [43, 44, 45, 41].
@@ IV Radiation Properties of the Thin Accretion Disk
where $\theta_{0}$ is the inclination angle of the observer, $\alpha$ is the polar angle of the photon’s position on the observer’s imaging plane, and $b$ is the impact parameter of the photon relative to the line joining the source and the observer [62, 41]. Using the LQRNBH metric, the above redshift factor can be written explicitly as
$1+z=\frac{1+\frac{r^{4}(r-Q^{2})-\zeta^{2}(r^{2}-2r+Q^{2})(r^{2}-4r+3Q^{2})}{(r^{2}-2r+Q^{2})(r^{4}+\zeta^{2}(r^{2}-2r+Q^{2}))}\sin\theta_{0}\sin\alpha}{\sqrt{\frac{(2Q^{2}+(-3+r)r)(r^{4}+2Q^{2}\zeta^{2}-4r\zeta^{2}+2r^{2}\zeta^{2})}{r^{6}}}}.$ (30)
@@ VI Conclusion and Discussion
The results of LQRNBH in this work provide several observational signals that can be used to distinguish it from the classical RNBH. However, our analysis focuses on the asymptotically flat LQRNBH and thin disk model. Subsequent research can be extended to the LQRNBH with nonzero cosmological constants and LQRNBH surrounded by other disk models [68, 69], or dynamical accretion disk [70]. It is also interesting to explore the features of the quasinormal modes of the LQRNBH. These further research may yield more insights on our understanding of the quantum gravity effects.

### v2

@@ Abstract
We investigate the characteristics of circular geodesics around loop quantum Reissner-Nordström black hole (LQRNBH) and the radiation properties and observational appearance of a thin accretion disk around it. By calculating the shadow radius and utilizing observational data from M87* and Sgr A*, we derive constraints on the quantum parameter $\zeta$ and charge parameter $Q$. The timelike circular geodesics around LQRNBH and the influence of the model parameters on the circular motion are also discussed. It is found that contrary to the case of the parameter $Q$, the increase of the quantum parameter $\zeta$ leads to the increase of the radius of the innermost stable circular orbit (ISCO). Then, by considering a thin accretion disk model, various radiation properties of the LQRNBH and the effects of the model parameters on them are studied. Concrete examples are provided for quantitative comparison of the radiation properties between LQRNBH and Schwarzschild black hole. With the ray-tracing method, the isoradial curves, redshift distributions, and the observed radiation fluxes of the direct and secondary images of the LQRNBH with the thin accretion disk for various model parameters and observation angles are numerically calculated and discussed. These results are beneficial for our understanding of the physical consequences of loop quantum gravity effect.
@@ I Introduction
The LQBHs have been studied extensively from various aspects. The quasinormal modes of these spherical LQBHs have been explored by several recent works [27, 28, 29, 30, 31, 32, 33, 34]. The correspondence between grey-body factors and quasinormal modes of these LQBHs were also studied [35]. The features of the shadows of these quantum black holes were studied in [36, 37, 38, 39, 40], and the constraints on the parameter encoding the quantum gravity effect from Event Horizon Telescope (EHT) observations [34, 41, 42, 43] and gravitational wave observation [44] were also discussed. Furthermore, the optical appearance of the accretion disk around LQBHs and the strong gravitational lensing effects of the LQBHs were explored in [45, 46, 47, 42]. The investigation of primordial LQBH accounts for the dark matter was studied in [48, 49].
@@ IV Radiation Properties of the Thin Accretion Disk
where $\theta_{0}$ is the inclination angle of the observer, $\alpha$ is the polar angle of the photon’s position on the observer’s imaging plane, and $b$ is the impact parameter of the photon relative to the line joining the black hole center and the observer [66, 42].
@@ VI Conclusion and Discussion
The results of LQRNBH in this work provide several observational signals that can be used to distinguish it from the classical RNBH. However, our analysis focuses on the asymptotically flat LQRNBH surrounded by a thin disk model. Subsequent research can be extended to consider LQRNBH surrounded by dark matter halo or thick disk models [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]. The polarization images and patterns of the accretion flow around a black hole, especially the near horizon one, can also be used to probe the black hole geometry [83, 84, 85, 86, 87], then to see how the loop quantum parameter affects these signatures and patterns can be future directions. It is also interesting to consider the LQRNBH with nonzero cosmological constant and explore the influence of model parameters on the dynamics of Einstein ring [88]. These further research may yield more insights on our understanding of the quantum gravity effects.